# What Does Neuro Mean to Cardio? Investigating the Role of Clinical Specialty Data in Medical LLMs

### v1

@@ Abstract
In this paper, we introduce S-MedQA, an English medical question-answering (QA) dataset for benchmarking large language models in fine-grained clinical specialties. We use S-MedQA to check the applicability of a popular hypothesis related to knowledge injection in the knowledge-intense scenario of medical QA, and show 1) that training on data from a speciality does not necessarily lead to best performance on that specialty and 2) regardless of the specialty fine-tuned on, token probabilities of clinically relevant terms for all specialties increase consistently. Thus, we believe improvement gains come mostly from domain shifting (e.g., general to medical) rather than knowledge injection and suggest rethinking the role of fine-tuning data in the medical domain. We release S-MedQA and all code needed to reproduce all our experiments to the research community. [^1] https://anonymous.4open.science/r/S-MedQA-85FD/
@@ 1 Introduction
We use our dataset to investigate a popular hypothesis that instruction tuning brings little-to-no knowledge into LLMs, as a few examples yield results comparable to large data (Zhou et al., 2024). Here, we define knowledge injection as the process of introducing new domain-specific information into a model during fine-tuning. To that regard, we first fine-tune LLMs on one specialty and evaluate on other specialties (see §3.2). Interestingly, the best results often come from fine-tuning on unrelated specialties. E.g., fine-tuning with infectious disease data performs best on cardiology domain, despite their knowledge being largely unrelated. Moreover, different pre-trained LLMs exhibits different knowledge transfer patterns across clinical specialties (see §A.12). In further experiments, we curate clinically relevant per-specialty terms (e.g., diseases, procedures) and analyze changes in these term probabilities before and after fine-tuning on data from different specialties. Our results suggest that performance gains are driven more by domain shifts (from general to medical domain) than by fine-grained clinical knowledge injection.
Figure 1: Overview of S-MedQA’s construction process. For single specialty annotation of each sample, we generate predictions using 5 different prompts and only keep those where predictions agree (3+limit-from33+3 +, 4+limit-from44+4 +, or $55$ times). We randomly sample 1,00010001,0001 , 000 questions from our train set and ask a medical expert to evaluate GPT-3.5’s predictions, achieving an accuracy ranging from 97.897.897.897.8% to 90.890.890.890.8%. Three medical students then annotate the same $200200$ samples out of the original 1,00010001,0001 , 000 annotated samples for computing inter-annotator agreements (see §2.4 for details). The expert also manually annotates S-MedQA’s whole validation and test set. For multi-specialty annotation, we leverage conformal prediction to assign labels across multiple clinical specialties, using a calibration set of $300300$ samples manually annotated by the medical expert and medical students. At a target confidence level of 90%, we determine a conformal threshold of $0.57390.5739$, which is then applied to all remaining samples for calibrated predictions.
@@ 2 A Benchmark of clinical specialties
We source examples from MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022), two widely used medical QA datasets. MedQA samples follow their original train/valid/test splits, whereas we only use MedMCQA’s training split for its test labels are not public.
@@ 2.1 Clinical specialty categorization
[…]
@@
Single specialty labels do not always capture the complexity of clinical questions. To account for this, we use conformal prediction (Angelopoulos and Bates, 2021) to generate calibrated multi-specialty labels in S-MedQA. At a 95% confidence level, we achieve an average coverage of 95% and an F1 score of 0.60 on the calibration set (see §A.5 for the methods and evaluation).
@@ 2.2 S-MedQA’s predicted specialties
In §A.3 (Fig. 3) we show the distribution of predicted specialties. We exclude $23102310$ (6666%) samples categorized as Others (containing clinically irrelevant information), and then focus on $1515$ out of $5555$ specialties with more than $300300$ samples to ensure statistical reliability. The final dataset comprises $4000740007$ / $899899$ / $893893$ samples in train/validation/test sets after human validation (described next in §2.4). For subsequent analyses, we use the top 6 specialties with the most samples (Cardiology, Gastroenterology, Infectious diseases, Neurology, Obstetrics and Gynecology, and Pediatrics).
@@ 2.4 Manual validation for single specialty predictions
A medical expert labels all the examples in S-MedQA’s validation and test sets with the single most correct clinical specialty. This expert also validates 1,00010001,0001 , 000 random samples from the train set, confirming whether the specialties predicted by GPT-3.5 are correct. [^4] This is the specialty that is the most relevant to the QA pair. In this evaluation, we exclude any cases where there is ambiguity in which clinical specialty is the most relevant. In general, when using voting with multiple prompts, we see large performance gains compared to using single prompts (e.g., from 72.872.872.872.880.280.280.280.2% to 90.890.890.890.897.897.897.897.8%; see §A.7).
@@
In Table 1, we show the accuracy vs. coverage trade-off over the 1,00010001,0001 , 000 random samples from the train set for different requirements for majority voting. A higher quorum results in higher accuracy ($90.8\rightarrow 97.890.8 → 97.8$) but greatly decreases the coverage ($89.1\rightarrow 49.289.1 → 49.2$). We release per-prompt single-specialty categorizations together with votes for all examples for users of the dataset to decide their preference between accuracy and coverage—more data but possibly more noise or less noise but less data—based on their specific use cases. We select ‘3+’ as the quorum in this study for adequate fine-tuning data.
Moreover, to assess the trustworthiness of the medical expert, we randomly sample $200200$ from the 1,00010001,0001 , 000 examples and further ask three medical graduate students to validate the same examples in the same procedure. We use Krippendorff’s alpha (Hayes and Krippendorff, 2007) to measure the inter-annotator agreement among the four annotators and obtain 83.6% (95% CI [69.0%, 93.9%]).
@@ 3.1 Experimental setup
We experiment on four variants of two open-source LLMs: Llama2-Chat-7b and 13b (Touvron et al., 2023) (Llama2-7b and Llama2-13b) and Mistral-Instruct-v0.1 and v0.2 (Jiang et al., 2023) (Mistral-v0.1 and Mistral-v0.2). We fine-tune each LLM on each of the six per-specialty training datasets with prompts shown in §A.9 and measure each resulting model’s performance on all six per-specialty test sets. In each experiment, we train for up to 10 epochs and perform model selection using per-specialty validation accuracy. We also train on the combined set containing all six specialties’ training data to evaluate how exposure to larger and more diverse data affects models’ performance. We provide more details and hyperparameters in §A.10.
We follow best practices to evaluate LLM performance on QA datasets by shuffling answers multiple times and adding multiple shuffled QA pairs in the test set, and by using the entire answer to a question instead of the LLM’s maximum token probability among options A, B, C, D in the answer (Zheng et al., 2023; Wang et al., 2024). Please see §A.11 for details on our evaluation protocol.
@@ 3.2 Results
Looking at the results of models fine-tuned on individual specialties, none of the best performing models were trained on the corresponding specialty’s training data, e.g., the best performance on the Cardio test set (55.4%) was achieved by the model trained on Infect. If the improvements were due to knowledge injection, we would expect to see best performing models consistently along the diagonal (e.g., Cardio $\rightarrow$ Cardio, Infect $\rightarrow$ Infect, and so on). We report results for other pretrained LLMs in §A.12 and note that similar conclusions hold whereas with different transfer patterns across specialties We note that Mistral-v0.1 behaves differently and shows a high within-specialty transfer; though we cannot explain the exact mechanism behind this behavior, we note that Mistral-v0.1 consistently underperforms v0.2 (in Table 2). We believe this requires further investigation and is an interesting avenue for future work.
@@ 4 Discussion
We hypothesize that the performance improvements in Table 2 are primarily due to domain shifting rather than knowledge injection; in other words, performance gains are due to domain adaptation from general LLMs to the clinical domain, rather than due to the injection of new clinical knowledge.
@@ A.5 Methods and evaluation of multi-specialty categorization
We first use GPT-3.5 to obtain a ranking of the predicted probabilities over all possible specialties for each sample. From this ranking, we apply conformal prediction to generate multi-specialty labels. Specifically, we use the calibration set to determine the threshold for including candidate specialties, so that the resulting predictions satisfy a predefined confidence level (95%). We dynamically adjust the number of specialties assigned to each sample, ensuring that the model captures the complexity of multi-specialty cases while maintaining statistical guarantees on the error rate. Therefore, only the top-ranked specialties meeting the threshold are included in the final label set.
We report key metrics computed on the calibration dataset to evaluate the performance of our multi-specialty categorization framework. The observed coverage rate, which measures the proportion of samples where the true labels are included in the predicted label set, was 95%, indicating a good calibration. The average number of predicted labels per sample was $2.462.46$, with an F-1 score of $0.600.60$.
@@ A.6 Clinical specialty benchmark description
In practice, we release all individual prompt predictions, as well as three versions of the dataset for majority voting with a minimum of 3+, 4+, and 5 votes. A coverage of 89.189.189.189.1% of the data leads to clinical specialties that are 90.890.890.890.8% accurate, whereas in the other side of the spectrum, we can obtain an accuracy of 97.897.897.897.8% while the coverage decreases to 49.249.249.249.2%. By sharing multiple versions of S-MedQA, we cater to different users’ needs. Users can then use more data (coverage of 89.189.189.189.1%) if their use-case can cope with mistakes in the order of 10101010% (majority voting 3+limit-from33+3 +); if the use-case requires data akin to gold-standard, i.e., error-free, users can use majority voting $55$ (which basically requires all 5 prompts to agree for an example to be included), which provides an accuracy of 97.897.897.897.8%.
@@
Table 4: Accuracy of each prompt. Prompts #$11$ to #$55$ are shown in Figures 59 in §A.2.
@@ A.11 LLM evaluation protocol
In all test sets, we shuffle the answers 5 times for each sample and add all these 5 entries to the final test set in case the model prefers an option due to its position (Zheng et al., 2023). To further improve the reliability of results, we follow Wang et al. (2024) to generate the entire answer with the model and train a classifier to match model outputs to the options in a post-hoc step, instead of using the maximum probability of options {A, B, C, D} with a single next-token prediction step. More concretely, we randomly select 150 training samples and generate answers for these with all four LLMs (Llama2-7b, Llama2-13b, Mistral-v0.1, and Mistral-v0.2), resulting in $600600$ responses. We manually annotate all the responses with the right options and use these annotations to train a Mistral-Instruct-v0.2 model as the classifier, with $400400$ ($200200$) train (test) samples. Our classifier achieves 96.5% accuracy and we use it in all experiments.
@@
A.12 Cross-specialty evaluation results for Mistral 7B v0.1, and Llama 7B and 13B
In Tables 67, and 8, we show cross-specialty evaluation matrices for Llama2-7b-chat, Llama2-13b-chat, and Mistral-7b-instruct-v0.1 in addition to our main results (in §3.2). Here we also observe that the best performance on each per-specialty test set is not achieved by the model that is tuned on training data from the same specialty for Llama models. With Mistral-7b-instruct-v0.1, however, the best performance is almost always obtained with the model trained on the same specialty.
When comparing these results with those obtained with Mistral-7b-instruct-v0.2 (in Table2 in our main paper), we note that results with Mistral-7b-instruct-v0.2 (in our main paper) are consistently better than those obtained with Mistral-7b-instruct-v0.1 (in Table 8). We believe that answering why some models better transfer within-specialty than across-specialties (or vice-versa) warrants further research on this topic.
Test Sets
Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
@@
Test Sets
Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
@@
Test Sets
Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.

### v2

@@ Abstract
In this paper, we introduce S-MedQA, an English medical question-answering (QA) dataset for benchmarking large language models in fine-grained clinical specialties. We use S-MedQA to check the applicability of a popular hypothesis related to knowledge injection in the knowledge-intense scenario of medical QA, and show that 1) training on data from a specialty does not necessarily lead to best performance on that specialty and 2) regardless of the specialty fine-tuned on, token probabilities of clinically relevant terms for all specialties increase consistently. Thus, we believe improvement gains come mostly from domain shifting (e.g., general to medical) rather than knowledge injection and suggest rethinking the role of fine-tuning data in the medical domain. We release S-MedQA and all code needed to reproduce all our experiments to the research community.  [^1] https://github.com/yanxinlan/S-MedQA
@@ 1 Introduction
We use our dataset to investigate the hypothesis that almost all knowledge in LLMs originates from pretraining, and that fine-tuning primarily serves to shift the model toward a specific knowledge domain (Zhou et al., 2024). To that regard, we first fine-tune LLMs on one specialty and evaluate on other specialties (see §3.2). Interestingly, the best results often come from fine-tuning on unrelated specialties. E.g., fine-tuning with infectious disease data performs best on cardiology domain, despite their knowledge being largely unrelated. Moreover, different pre-trained LLMs exhibits different knowledge transfer patterns across clinical specialties (see §A.12). In further experiments, we curate clinically relevant per-specialty terms (e.g., diseases, procedures) and analyze changes in these term probabilities before and after fine-tuning on data from different specialties. Our results suggest that performance gains are driven more by domain shifts (from general to medical domain) than by fine-grained clinical knowledge injection.
Figure 1: Overview of S-MedQA’s construction process. For single specialty annotation of each sample, we generate predictions using 5 different prompts and only keep those where predictions agree ($3+$, $4+$, or $5$ times). We randomly sample $1,000$ questions from our train set and ask a medical expert to evaluate GPT-3.5’s predictions, achieving an accuracy ranging from $97.8$% to $90.8$%. Three medical students then annotate the same $200$ samples out of the original $1,000$ annotated samples for computing inter-annotator agreements (see §2.4 for details). The expert also manually annotates S-MedQA’s whole validation and test set. For multi-specialty annotation, we leverage conformal prediction to assign labels across multiple clinical specialties, using a calibration set of $300$ samples manually annotated by the medical expert and medical students. At a target confidence level of 90%, we determine a conformal threshold of $0.5739$, which is then applied to all remaining samples for calibrated predictions.
@@ 2 A Benchmark of clinical specialties
We source examples from MedQA (Jin et al., 2021) and MedMCQA (Pal et al., 2022), two widely used medical QA datasets. MedQA samples follow their original train/valid/test splits, whereas we only use MedMCQA’s training split as its test labels are not public.
@@ 2.1 Clinical specialty categorization
Manually labeling QA examples with clinical specialties is costly and time-consuming. To reduce manual effort, we use GPT-3.5 to annotate samples with single clinical specialties. However, preliminary experiments with a single prompt to predict a single specialty achieve low accuracy($\sim 75$%). To improve this, we design five prompts (details in §A.2), generate single specialty predictions with GPT-3.5 for each Q&A sample, and apply majority voting to select the specialty (Ding et al., 2023; Goel et al., 2023). Human evaluation shows accuracies between $90.8$%–$97.8$% (see §2.4 for details).
@@
Single specialty labels do not always capture the complexity of clinical questions. To account for this, we applied conformal prediction (Angelopoulos and Bates, 2021) to generate calibrated multi-specialty labels in S-MedQA. At a 95% confidence level, we achieve an average coverage of 95% and an F1 score of 0.60 on the calibration set (see §A.5 for the methods and evaluation).
@@ 2.2 S-MedQA’s predicted specialties
In §A.3 (Fig. 3) we show the distribution of predicted specialties. We exclude $2310$ ($6$%) samples categorized as Others (containing clinically irrelevant information), and then focus on $15$ out of $55$ specialties with more than $300$ samples to ensure statistical reliability. The final dataset comprises $40007$ / $899$ / $893$ samples in train/validation/test sets after human validation (described next in §2.4). For subsequent analyses, we use the top 6 specialties with the most samples (Cardiology, Gastroenterology, Infectious diseases, Neurology, Obstetrics and Gynecology, and Pediatrics).
@@ 2.4 Manual validation for single specialty predictions
A medical expert labels all the examples in S-MedQA’s validation and test sets with the single most correct clinical specialty. This expert also validates $1,000$ random samples from the train set, confirming whether the specialties predicted by GPT-3.5 are correct. [^4] This is the specialty that is the most relevant to the QA pair. In this evaluation, we exclude any cases where there is ambiguity in which clinical specialty is the most relevant. In general, when using voting with multiple prompts, we see large performance gains compared to using single prompts (e.g., from 72.872.872.872.880.280.280.280.2% to 90.890.890.890.897.897.897.897.8%; see §A.7).
@@
In Table 1, we show the accuracy vs. coverage trade-off over the $1,000$ random samples from the train set for different requirements for majority voting. A higher quorum results in higher accuracy ($90.8\rightarrow 97.8$) but greatly decreases the coverage ($89.1\rightarrow 49.2$). We release per-prompt single-specialty categorizations together with votes for all examples for users of the dataset to decide their preference between accuracy and coverage—more data but possibly more noise or less noise but less data—based on their specific use cases. We select ‘3+’ as the quorum in this study for adequate fine-tuning data.
Moreover, to assess the trustworthiness of the medical expert, we randomly sample $200$ from the $1,000$ examples and further ask three medical graduate students to validate the same examples in the same procedure. We use Krippendorff’s alpha (Hayes and Krippendorff, 2007) to measure the inter-annotator agreement among the four annotators and obtain 83.6% (95% CI [69.0%, 93.9%]).
@@ 3.1 Experimental setup
We experiment on 8 open-source LLMs. Six general-purpose models: Llama2-Chat-7B, Llama2 -Chat-13B, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (Touvron et al., 2023), Mistral-Instruct-v0.1 and v0.2 (Jiang et al., 2023), one biomedical model: Bio-Medical-LLaMA-3-8B. We also include OLMo-2-1124-7B-Instruct to rule out the possible contamination of training data.
We fine-tune each LLM on the six per-specialty training sets using prompts from §A.9, and evaluate on all six test sets. Models are trained for up to 10 epochs, with selection based on per-specialty validation accuracy. We also train on the combined dataset to evaluate how exposure to larger and more diverse data affects models’ performance. Additional details and hyperparameters are in §A.10.
We follow best practices to evaluate LLM performance on QA datasets by shuffling answers multiple times and adding multiple shuffled QA pairs in the test set, and by using the entire answer to a question instead of the LLM’s maximum token probability among options A, B, C, D in the answer (Zheng et al., 2023; Wang et al., 2024). Please see §A.11 for details on our evaluation protocol. To verify the effectiveness of the fine-tuning process, we assess the model’s performance on the training data and achieve an average accuracy of 89.3% across the fine-tuning datasets.
@@ 3.2 Results
Looking at the results of models fine-tuned on individual specialties, none of the best performing models were trained on the corresponding specialty’s training data, e.g., the best performance on the Cardio test set (55.4%) was achieved by the model trained on Infect. If the improvements were due to knowledge injection, we would expect to see best performing models consistently along the diagonal (e.g., Cardio $\rightarrow$ Cardio, Infect $\rightarrow$ Infect, and so on). We report results for other pretrained LLMs in §A.12 and note that similar conclusions hold whereas with different transfer patterns across specialties.
@@ 4 Discussion
We hypothesize that the performance improvements in Table 2 are primarily due to domain shifting from general LLMs to the clinical domain, rather than due to the injection of new clinical knowledge.
@@
We further fine-tune Mistral-v0.2 on 3 Social Sciences subsets of MMLU (public relations, security studies, and sociology). There remains a gap between the average log probabilities of all six specialty terms generated by the model compared to the model trained on S-MedQA, ensuring that the observed probability increase is a result of domain-specific adaptation rather than a mere consequence of additional training (See §A.13 for detail).
@@ 5 Conclusions
Acknowledgment
XY and IC are funded by the project CaRe-NLP with file number NGF.1607.22.014 of the research programme AiNed Fellowship Grants which is (partly) financed by the Dutch Research Council (NWO).
@@ A.5 Methods and evaluation of multi-specialty categorization
For each example in the calibration set, we compute log-probabilities for all 55 candidate specialties using GPT-3.5. This is done iteratively: at each step, we extract the top-k (with k=20) next-token predictions from the model, and extend the input with tokens that could lead to valid specialty names. This process continue until either all 55 specialties were generated — allowing us to collect their corresponding token-level log-probabilities — or the next token required for a specialty does not appear in the top-20 predictions, in which case we assign that specialty a probability of zero. For each successfully generated specialty, we compute a score as the average of its next-token log-probabilities, and normalize these scores to obtain a probability distribution over the 55 specialties. We then define the conformal score as $1-P$, where $P$ is the normalized probability of a given specialty, and use this to construct prediction sets with a 95% confidence level ($\alpha$=0.05). Since each question may have multiple correct specialties, we select $P$ as the highest predicted probability among the set of annotated correct specialties for evaluation.
We report key metrics computed on the calibration dataset to evaluate the performance of our multi-specialty categorization framework. The observed coverage rate, which measures the proportion of samples where the true labels are included in the predicted label set, was 95%, indicating a good calibration. The average number of predicted labels per sample was $2.46$, with an F-1 score of $0.60$.
@@ A.6 Clinical specialty benchmark description
In practice, we release all individual prompt predictions, as well as three versions of the dataset for majority voting with a minimum of 3+, 4+, and 5 votes. A coverage of $89.1$% of the data leads to clinical specialties that are $90.8$% accurate, whereas in the other side of the spectrum, we can obtain an accuracy of $97.8$% while the coverage decreases to $49.2$%. By sharing multiple versions of S-MedQA, we cater to different users’ needs. Users can then use more data (coverage of $89.1$%) if their use-case can cope with mistakes in the order of $10$% (majority voting $3+$); if the use-case requires data akin to gold-standard, i.e., error-free, users can use majority voting $5$ (which basically requires all 5 prompts to agree for an example to be included), which provides an accuracy of $97.8$%.
@@
Table 4: Accuracy of each prompt. Prompts #$1$ to #$5$ are shown in Figures 59 in §A.2.
@@ A.11 LLM evaluation protocol
In all test sets, we shuffle the answers 5 times for each sample and add all these 5 entries to the final test set in case the model prefers an option due to its position (Zheng et al., 2023). To further improve the reliability of results, we follow Wang et al. (2024) to generate the entire answer with the model and train a classifier to match model outputs to the options in a post-hoc step, instead of using the maximum probability of options {A, B, C, D} with a single next-token prediction step. More concretely, we randomly select 150 training samples and generate answers for these with all four LLMs (Llama2-7b, Llama2-13b, Mistral-v0.1, and Mistral-v0.2), resulting in $600$ responses. We manually annotate all the responses with the right options and use these annotations to train a Mistral-Instruct-v0.2 model as the classifier, with $400$ ($200$) train (test) samples. Our classifier achieves 96.5% accuracy and we use it in all experiments.
@@
A.12 Cross-specialty evaluation results for Mistral-7B-v0.1, Llama-2-7B and -13B, Llama-3.1-8B, Llama-3.2-3B, Bio-Medical-LLaMA-3-8B, and OLMo-2-1124-7B-Instruct
In Tables 78,  9,  10,  11,  12,  13 we show cross-specialty evaluation matrices for Llama2-7b-chat, Llama2-13b-chat, Mistral-7b-instruct-v0.1, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Bio-Medical-LLaMA-3-8B, and and OLMo-2-1124-7B-Instruct in addition to our main results (in §3.2). Here we also observe that the best performance on each per-specialty test set is not achieved by the model that is tuned on training data from the same specialty for Llama models.
A.13 Average token log probabilities gap between model fine-tuned on Social Science subsets vs. S-MedQA
We fine-tune Mistral-v0.2 on three Social Sciences subsets of MMLU—public relations, security studies, and sociology—which are entirely unrelated to the medical domain. We compare the average token log-probabilities of six specialty medical terms produced by this model with those obtained from models fine-tuned on the corresponding S-MedQA datasets. As shown in Table 6, there is a significant drop in token probabilities when fine-tuned on out-of-domain data, confirming that the change in token probabilities is brought by domain shifting rather than only addition training.
Table 6: Token probability increase by clinical specialty after fine-tuning.
Specialty Increase (%)
Cardio 10.5
Gastro 10.7
Infect 9.5
Neuro 8.4
Obstetrics 7.2
Pediatrics 7.5
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
(Sample size) (80) (83) (102) (74) (88) (90)
Llama2-7b 36.0 36.3 36.7 34.6 40.6 41.4 37.7
@@
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
(Sample size) (80) (83) (102) (74) (88) (90)
Llama2-13b 43.3 34.9 40.6 36.1 45.4 39.2 40.0
@@
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
(Sample size) (80) (83) (102) (74) (88) (90)
Mistral-v0.1 41.0 39.0 40.0 30.7 42.7 37.5 38.9
@@
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
(Sample size) (80) (83) (102) (74) (88) (90)
Train Sets Llama-3.1 64.0 67.5 65.3 66.8 65.0 64.5 65.5
Cardio 69.3 76.5 73.5 76.2 70.5 70.9 72.8
Gastro 68.2 73.3 71.8 74.6 69.3 69.2 71.1
Infect 65.8 74.2 69.5 72.5 66.6 66.7 69.3
Neuro 69.9 76.4 72.3 76.9 69.5 71.7 72.8
Obstetrics 69.8 75.0 72.9 75.4 68.7 69.4 71.9
Pediatrics 66.9 75.0 72.1 75.9 68.7 70.2 71.4
Combined 72.9 79.5 75.4 77.8 70.8 74.1 75.1
Table 10: Cross-specialty accuracy matrix of Llama-3.1-8B-Instruct.
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
(Sample size) (80) (83) (102) (74) (88) (90)
Train Sets Llama-3.2 54.0 60.0 58.5 61.0 57.0 55.5 57.7
Cardio 59.8 68.2 67.9 70.6 62.5 59.0 64.7
Gastro 58.0 65.3 66.2 68.7 60.2 56.5 62.5
Infect 59.5 63.9 65.7 69.0 62.5 57.8 63.1
Neuro 59.8 64.6 66.9 69.8 62.2 58.6 63.6
Obstetrics 60.1 65.6 67.1 70.2 61.6 58.9 63.9
Pediatrics 61.0 66.3 67.7 70.3 62.0 58.6 64.3
Combined 65.4 69.9 71.3 75.5 66.9 65.0 68.9
Table 11: Cross-specialty accuracy matrix of Llama-3.2-3B-Instruct.
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
(Sample size) (80) (83) (102) (74) (88) (90)
Train Sets Bio-Llama 70.1 72.3 71.0 72.8 70.2 69.4 71.0
Cardio 79.7 82.7 80.3 82.8 76.4 75.3 79.6
Gastro 77.7 83.1 81.5 82.8 76.6 74.9 79.5
Infect 76.8 82.2 78.9 82.2 75.1 72.8 78.1
Neuro 77.1 80.7 79.7 83.8 76.3 74.0 78.6
Obstetrics 78.0 81.7 81.2 81.8 76.6 73.7 78.9
Pediatrics 79.3 82.3 81.3 82.8 78.2 75.3 79.9
Combined 82.7 84.3 82.0 85.5 79.6 79.5 82.3
Table 12: Cross-specialty accuracy matrix of Bio-Medical-Llama-3-8B.
Test Sets Cardio Gastro Infect Neuro Obstetrics Pediatrics avg.
(Sample size) (80) (83) (102) (74) (88) (90)
Train Sets OLMo 36.0 39.5 40.3 38.8 37.2 35.9 38.0
Cardio 40.7 46.0 47.2 44.4 41.6 38.2 43.1
Gastro 39.6 46.4 47.7 45.7 41.0 35.7 42.8
Infect 38.4 45.5 49.0 45.7 40.2 37.2 42.7
Neuro 39.0 46.4 46.5 45.4 40.7 38.4 42.8
Obstetrics 41.5 44.5 46.1 45.3 39.5 37.6 42.4
Pediatrics 39.6 46.9 48.5 43.5 39.7 37.8 42.7
Combined 48.7 54.1 55.6 55.5 48.1 42.2 50.8
Table 13: Cross-specialty accuracy matrix of OLMo-2-1124-7B-Instruct.